OpenStudy (anonymous):

solve the solution then classify the system as consistent and independent ,consistent and dependent or inconsistent, then graph x-2y=4 4x+2y=6 *list the points so i can graph

OpenStudy (anonymous):

x-2y=4, (*-4) -4x+8y=-16 4x+2y=6 _____________ 10y=-10, y=-1 x-2(-1)=4 x+2=4, x=2. A(2,-1) consistent, independent 

OpenStudy (saifoo.khan):

x = 2..................y = -1

OpenStudy (anonymous):

okay thanks but what are the points to graph it on my own ?

OpenStudy (anonymous):

...

OpenStudy (saifoo.khan):

u can graph that's easy!

OpenStudy (anonymous):

i know i just need 3 ponts to graph it

OpenStudy (anonymous):

-1 0 1